The human heart forces 4500 cc of blood per min through the arteries under a pressure of 15 cm of hg.calculate the hp of the heart. Take 1 hp=746 w and g =980 cm/second square.

Grade:9

2 Answers

Himalaya Nirjhar Datta
7 years ago
I think few data is missing.Like it is pumping up or down etc. More over densities are also not given. Power is work/timeWork is force/displacement.Displacement is actually 0 in a cycle.Hence work is 0.So, power is also 0 Watt= 0 HP
Xyz
6 years ago
First we need to convert 15 cm of Hg into Si unts . 15cm of Hg =19998.36 kg/ms^2Then convert volume of flow in Si 4500cm^3/min = 0.000075m^3/secUse the formula Power =PRESSURE ×VOL. OF FLOW PER SECP=(19998.36 × .000075) = 1.5WCONVERT INTO HPPOWER =( 1.5/746)= 20 × 10^-4 HP (APPROX.